a titration of a 25.00 ml sample of a hydrochloric acid solution of unknown molarity reaches the equivalence point when 39.12 ml of 0.3204 M naOh solution has been added. what is the molarity of the Hcl solution



1
Expert's answer
2021-07-27T11:19:04-0400

NaOH + HCl = NaCl + H2O

According to the equation, n (NaOH) = n (HCl).

CM = n / V

n = CM x V

CM (NaOH) x V (NaOH) = CM (HCl) x V (HCl)

0.3204 x 0.03912 = CM (HCl) x 0.025

CM (HCl) = (0.3204 x 0.03912) / 0.025 = 0.5 M